. Suppose the Demand for coffee is given by: Q=25-3p and the current market price is $3. Find the Consumer Surplus (CS). (3pts
Answer #1

Answer

the demand function is:

Q=25-3P

P=3

Q=25-3*3=16

--------

Converting the demand function to its inverse

Q=25-3P

3P=25-Q

P=(25/3)-(1/3)Q

CS is the area above price and below the demand curve

CS=0.5*( Y-axis intercept of the demand curve -P)*Q

=0.5*((25/3)-3)*16

=$42.6666667

the CS is $42.67